# Weakly Supervised Concept Learning with Class-Level Priors for Interpretable Medical Diagnosis

### v1

@@ 1 Introduction
Deep learning (DL) has achieved remarkable success in medical imaging, yet most models operate as “black boxes”, limiting clinical trust and adoption. Interpretable-by-design (IBD) models, such as Concept Bottleneck Models (CBM) [1] and Variational-Information Pursuit (V-IP) [2], address this by mapping raw image features to human-understandable concepts (e.g., blue-whitish veil, irregular streaks) and using them for classification. CBMs require all concepts to be predicted before making a decision, while V-IP adaptively queries only the most informative ones. Both approaches provide meaningful explanations aligned with clinical reasoning. Building on these ideas, the Multimodal Interpretable Concept Alignment (MICA) [3] and Concept-Based Vision–Language Model (CBVLM) [4] require few-shot or partially annotated concepts. However, collecting per-concept annotations at scale remains difficult in medical imaging, as concepts are subtle (there is often disagreement even amongst experts) and experts have limited time and resources. Thus, the central challenge is achieving concept prediction without explicit concept supervision.
Recent works attempt to relax this dependency. Post-hoc CBMs [5] use Concept Activation Vectors from external datasets but fail when no curated concept bank exists. Language-guided models such as LaBo-CBM [6] use large language models (LLMs) to define textual concepts, while V2C-CBM [7] employs VLMs to tokenize visual concepts. Similarly, ConceptCLIP [8] learns image–text–concept embeddings via medical ontologies, but its reliance on predefined vocabularies and alignment quality limits generalization to subtle clinical findings. Recent V-IP extensions, including Bootstrapped [9] and Learned-Query V-IP [10], integrate VLMs like CLIP to answer or learn interpretable queries but still rely on textual supervision. In a different line of work, the Visual Perception Module (VPM) [11] introduces attention-based refinement for concept-guided zero-shot classification but assumes access to ground-truth concept annotations, differing from our goal of predicting concept answers without annotations or VLM reliance.
@@ 2.2 Proposed Model
We use element-wise multiplication to combine the projected image features with the surrogate concept vectors, akin to VPM [11]:
@@
The attention vector $\gamma(x)$ highlights the concepts most relevant for predicting class $y$. Directly masking with $\gamma(x)$ as in [11] risks discarding weak but potentially informative concepts. To mitigate this, we adopt a residual refinement mechanism:
@@
where $\beta$ is a tunable parameter controlling the strength of refinement. This ensures that important concepts are amplified while less relevant but possibly useful ones are not completely suppressed, improving both stability and robustness.
@@ 2.3 Training Objective
(1) Triplet Loss. Following [11], we apply a triplet loss on the refined embeddings $\mathbf{z}^{\prime}$. For each anchor $\mathbf{z}^{\prime}_{i}$, we randomly select a positive example $\mathbf{z}^{\prime}_{j}$ from the same class and a negative example $\mathbf{z}^{\prime}_{m}$ from a different class. The triplet loss is defined as
@@ 3 Experiments
We evaluate the proposed framework on four medical imaging datasets: PH2 [12], a small binary dermoscopy dataset with 200 images and 8 clinical concepts; WBCatt [13], a large-scale hematology dataset with 10,198 images from five white blood cell types and 29 morphological concepts; HAM10000 [14], a dermoscopic dataset with 7,818 images from two diagnostic classes (nevus and melanoma) and 25 concepts; and CXR4 [15, 16], a chest X-ray dataset with 7,135 images from four diagnostic classes (Normal, COVID-19, Pneumonia, and Tuberculosis) and 29 radiological concepts. For the HAM10000 and CXR4 datasets, concept priors were generated using an LLM (e.g., ChatGPT) based on clinically meaningful class descriptions.
@@ 3.1 Results and Discussion
All experiments were conducted with three random seeds (except for CLIP-based models, which are deterministic) and train-validation-test splits to ensure robustness. We evaluate both concept prediction and downstream classification performance on WBCatt and PH2, where concept-level annotations are available, to verify that our method produces interpretable and predictive concept representations. In our setup, Baseline refers to models trained with ground-truth concepts; Vanilla-CBM/V-IP denote the standard CBM [1] and V-IP [2] trained with full concept supervision; and PCP-CBM and PCP-V-IP correspond to our proposed weakly supervised models trained only with class-level priors.
@@
CLIP [17] $41.25$ $41.41$ $56.56$ $32.96$
SigLIP [18] $58.05$ $4.58$ $61.25$ $24.79$
BioMedCLIP [19] $55.52$ $12.21$ $61.56$ $43.81$
ConceptCLIP [8] $48.43$ $26.34$ $58.75$ $39.33$
@@
Concept Prediction. In Table 1 we compare against zero-shot VLM-based methods (CLIP [17], SigLip [18], BioMedCLIP [19], ConceptCLIP [8]). As is seen, these methods perform poorly in medical imaging, likely caused by a lack of medical data seen during pretraining. However, by learning from minimal extra data in the from of concept-class priors, our prior-guided PCP framework yields stronger and more clinically consistent predictions with concept-level accuracy improvements of more than 30% compared to zero- shot baselines.
@@
Vanilla-CBM [1] $96.11_{\pm 0.24}$ $77.14_{\pm 0.02}$ - -
Vanilla-V-IP [2] $96.31_{\pm 0.12}$ $90.00_{\pm 0.78}$ - -
@@
For HAM10000 and CXR4, where concept annotations are unavailable (Vanilla-CBM/V-IP require concept annotations), we compare our models against the BlackBox (ResNet as backbone) baseline. Both PCP-CBM and PCP-V-IP achieve competitive F1-scores, demonstrating that our framework can learn meaningful concept representations directly from class-level priors, even in the absence of explicit concept supervision.
@@ 4 Conclusion
In this work, we present PCP, a weakly supervised framework that enables medical concept prediction without explicit concept annotations or reliance on vision–language supervision. By leveraging class-level concept priors and regularizing with KL and entropy losses, PCP achieves interpretable and clinically aligned concept predictions while maintaining competitive classification performance across multiple modalities. Although PCP utilizes class-level priors, these are far easier to obtain than exhaustive concept annotations and can be estimated from expert knowledge or dataset statistics, making the approach practical and scalable. Future work will focus on adaptive prior refinement and self-distilled concept reasoning to further enhance generalizability and reduce reliance on fixed priors.

### v2

@@ 1 Introduction
Deep learning (DL) has achieved remarkable success in medical imaging, yet most models operate as “black boxes”, limiting clinical trust and adoption. Interpretable-by-design (IBD) models, such as Concept Bottleneck Models (CBM) [koh2020concept] and Variational-Information Pursuit (V-IP) [chattopadhyayvariational], address this by mapping raw image features to human-understandable concepts (e.g., blue-whitish veil, irregular streaks) and using them for classification. CBMs require all concepts to be predicted before making a decision, while V-IP adaptively queries only the most informative ones. Both approaches provide meaningful explanations aligned with clinical reasoning. Building on these ideas, the Multimodal Interpretable Concept Alignment (MICA) [bie2024mica] and Concept-Based Vision–Language Model (CBVLM) [patricio2025cbvlm] require few-shot or partially annotated concepts. However, collecting per-concept annotations at scale remains difficult in medical imaging, as concepts are subtle (there is often disagreement even amongst experts) and experts have limited time and resources. Thus, the central challenge is achieving concept prediction without explicit concept supervision.
Recent works attempt to relax this dependency. Post-hoc CBMs [yuksekgonulpost] use Concept Activation Vectors from external datasets but fail when no curated concept bank exists. Language-guided models such as LaBo-CBM [yang2023language] use large language models (LLMs) to define textual concepts, while V2C-CBM [he2025v2c] employs VLMs to tokenize visual concepts. Similarly, ConceptCLIP [nie2025conceptclip] learns image–text–concept embeddings via medical ontologies, but its reliance on predefined vocabularies and alignment quality limits generalization to subtle clinical findings. Recent V-IP extensions, including Bootstrapped [chattopadhyay2024bootstrapping] and Learned-Query V-IP [kolek2025learning], integrate VLMs like CLIP to answer or learn interpretable queries but still rely on textual supervision. In a different line of work, the Visual Perception Module (VPM) [yu2024zero] introduces attention-based refinement for concept-guided zero-shot classification but assumes access to ground-truth concept annotations, differing from our goal of predicting concept answers without annotations or VLM reliance.
@@ 2.2 Proposed Model
We use element-wise multiplication to combine the projected image features with the surrogate concept vectors, akin to VPM [yu2024zero]:
@@
The attention vector $\gamma(x)$ highlights the concepts most relevant for predicting class $y$. Directly masking with $\gamma(x)$ as in [yu2024zero] risks discarding weak but potentially informative concepts. To mitigate this, we adopt a residual refinement mechanism:
@@
where $\beta$ is a tunable refinement strength parameter selected on the validation set for each dataset, amplifying important concepts while avoiding complete suppression of weaker but potentially informative ones, thereby improving stability and robustness.
@@ 2.3 Training Objective
(1) Triplet Loss. Following [yu2024zero], we apply a triplet loss on the refined embeddings $\mathbf{z}^{\prime}$. For each anchor $\mathbf{z}^{\prime}_{i}$, we randomly select a positive example $\mathbf{z}^{\prime}_{j}$ from the same class and a negative example $\mathbf{z}^{\prime}_{m}$ from a different class. The triplet loss is defined as
@@ 3 Experiments
We evaluate the proposed framework on four medical imaging datasets: PH2 [mendoncca2013ph], a small binary dermoscopy dataset with 200 images and 8 clinical concepts; WBCatt [tsutsui2023wbcatt], a large-scale hematology dataset with 10,198 images from five white blood cell types and 29 morphological concepts; HAM10000 [tschandl2018ham10000], a dermoscopic dataset with 7,818 images from two diagnostic classes (nevus and melanoma) and 25 concepts; and CXR4 [T.rahman2020reliable, cohen2020covidProspective], a chest X-ray dataset with 7,135 images from four diagnostic classes (Normal, COVID-19, Pneumonia, and Tuberculosis) and 29 radiological concepts. For the HAM10000 and CXR4 datasets, concept priors were generated using an LLM (e.g., ChatGPT) based on clinically meaningful class descriptions.
@@ 3.1 Results and Discussion
All experiments were conducted with three random seeds (except for CLIP-based models, which are deterministic) and train-validation-test splits to ensure robustness. We evaluate both concept prediction and downstream classification performance on WBCatt and PH2, where concept-level annotations are available, to verify that our method produces interpretable and predictive concept representations. In our setup, Baseline refers to models trained with ground-truth concepts; Vanilla-CBM/V-IP denote the standard CBM [koh2020concept] and V-IP [chattopadhyayvariational] trained with full concept supervision; and PCP-CBM and PCP-V-IP correspond to our proposed weakly supervised models trained only with class-level priors.
@@
CLIP [radford2021learning] $41.25$ $41.41$ $56.56$ $32.96$
SigLIP [zhai2023sigmoid] $58.05$ $4.58$ $61.25$ $24.79$
BioMedCLIP [zhang2023biomedclip] $55.52$ $12.21$ $61.56$ $43.81$
ConceptCLIP [nie2025conceptclip] $48.43$ $26.34$ $58.75$ $39.33$
@@
Concept Prediction. In Table 1 we compare against zero-shot VLM-based methods (CLIP [radford2021learning], SigLip [zhai2023sigmoid], BioMedCLIP [zhang2023biomedclip], ConceptCLIP [nie2025conceptclip]). As is seen, these methods perform poorly in medical imaging, likely caused by a lack of medical data seen during pretraining. However, by learning from minimal extra data in the from of concept-class priors, our prior-guided PCP framework yields stronger and more clinically consistent predictions with concept-level accuracy improvements of more than 30% compared to zero- shot baselines.
@@
Vanilla-CBM [koh2020concept] $96.11_{\pm 0.24}$ $77.14_{\pm 0.02}$ - -
Vanilla-V-IP [chattopadhyayvariational] $96.31_{\pm 0.12}$ $90.00_{\pm 0.78}$ - -
@@
For HAM10000 and CXR4, which lack concept annotations (required by Vanilla-CBM/V-IP), we compare against a ResNet black-box baseline. Both PCP-CBM and PCP-V-IP achieve competitive F1-scores, indicating that meaningful concept representations can be learned directly from class-level priors without explicit concept supervision.
Figure 2: Qualitative interpretability examples using PCP. Left: CBM-based prediction on PH2. Right: V-IP-based prediction on WBCatt.
Qualitative Analysis. Fig. 2 illustrates qualitative interpretability results on dermoscopy (PH2) and hematology (WBCatt). The model predicts clinically relevant concepts with high confidence, producing interpretable reasoning patterns consistent with expert knowledge, despite being trained without concept-level annotations.
@@ 4 Conclusion
In this work, we present PCP, a weakly supervised framework that enables medical concept prediction without explicit concept annotations or reliance on vision–language supervision. By leveraging class-level concept priors with KL and entropy regularization, PCP produces interpretable and clinically aligned concept predictions while maintaining competitive classification performance across multiple modalities. Although PCP relies on class-level priors, these are substantially easier to obtain than exhaustive concept annotations and can be derived from expert knowledge or dataset statistics, making the approach practical and scalable. In rare-disease or limited-data settings, prior quality may be noisier, which we identify as a key limitation and motivation for future work on adaptive prior refinement and self-distilled concept reasoning to improve robustness and generalization.